OpenStudy (anonymous):

how to factor 7y^3-448?

OpenStudy (anonymous):

The only thing you can factor here is the 7 (if I'm reading your equation correctly: \[7((y ^{3})-64)\]

OpenStudy (shamim):

7y^3-448 =7(y^3-64) =7(y^3-4^3) =7(y-4)(y^2+4y+4^2) =7(y-4)(y^2+4y+16)